Write the equivalent of these

quantities in SI unit with scientific

notation by showing step by step

conversion.


a) 0.30m = ______ mm

b) 5.00 mm = _____ cm

c) 0.400 L = _______ mL

d) 345 g = ______ kg


1
Expert's answer
2021-08-18T02:24:29-0400

a) 0.30m is 3·10-1·1·103 = 3·102 mm. SI: 3·10-1 m.

b) 5.00 mm is 5·1·10-2 cm = 5·10-2 cm. SI: 5·10-3 m.

c) 0.400 L is 0.4·1·103 = 0.4·103 mL. SI: 4·10-4 m3.

d) 345 g is 345·1·10-3 = 345·10-3 kg. SI: 3.45·102 kg.